\begin{thebibliography}{10}

\bibitem{key-1}
S.~Abbasbandy, S.~Kazem, M.~S. Alhuthali, and H.~H. Alsulami, \emph{Application
  of the operational matrix of fractional-order {L}egendre functions for
  solving the time-fractional convection-diffusion equation}, Appl. Math.
  Comput. \textbf{266} (2015), 31--40. \MR{3377539}

\bibitem{key-2}
M.~Abramowitz and I.~A. Stegun, \emph{Handbook of mathematical functions with
  formulas, graphs, and mathematical tables}, National Bureau of Standards
  Applied Mathematics Series, vol.~55, For sale by the Superintendent of
  Documents, U.S. Government Printing Office, Washington, D.C., 1964.
  \MR{0167642 (29 \#4914)}

\bibitem{key-3}
G.~E. Albert, \emph{Asymtotic forms for a general class of hypergeometric
  functions with aapplication to the generalized {L}egendre functions}, Duke.
  Math. J. \textbf{5} (1939), no.~2, 281--292.

\bibitem{key-4}
G.~B. Arfken and H.~J. Weber, \emph{Mathematical methods for physicists}, fifth
  ed., Harcourt/Academic Press, Burlington, MA, 2001. \MR{1810939
  (2001j:00004)}

\bibitem{key-5}
V.~M. Babich, \emph{To the problem on the asymptotics with respect to the
  indices of the associated {L}egendre function}, Russ. J. Math. Phys.
  \textbf{9} (2002), no.~1, 6--13. \MR{1965504 (2004c:33006)}

\bibitem{key-6}
W.~N. Bailey, \emph{On the {P}roduct of two {L}egendre {P}olynomials with
  {D}ifferent {A}rguments}, Proc. London Math. Soc. \textbf{S2-41}, no.~3, 215.
  \MR{1577113}

\bibitem{key-7}
{\c{S}}.~S. Bayin, \emph{Mathematical methods in science and engineering},
  Wiley-Interscience [John Wiley \& Sons], Hoboken, NJ, 2006. \MR{2238528
  (2007a:00001)}

\bibitem{key-8}
D.~Beckwith, \emph{Notes: {L}egendre {P}olynomials and {P}olygon
  {D}issections?}, Amer. Math. Monthly \textbf{105} (1998), no.~3, 256--257.
  \MR{1543182}

\bibitem{key-9}
S.~L. Belousov, \emph{Tables of normalized associated {L}egendre polynomials},
  Mathematical tables, vol.~18, Pergamon Press, 1962.

\bibitem{key-10}
W.~C. Brenke, \emph{On the summability and generalized sum of a series of
  {L}egendre polynomials}, Bull. Amer. Math. Soc. \textbf{39} (1933), no.~10,
  821--824. \MR{1562735}

\bibitem{key-15}
L.~Carlitz, \emph{Some arithmetic properties of the {L}egendre polynomials},
  Acta Arith. \textbf{4} (1958), 99--107. \MR{0096828 (20 \#3310)}

\bibitem{key-16}
\bysame, \emph{Some congruence properties of the {L}egendre polynomials}, Math.
  Mag. \textbf{34} (1960/1961), 387--390. \MR{0138799 (25 \#2242)}

\bibitem{key-14}
\bysame, \emph{Some integrals containing products of {L}egendre polynomials},
  Arch. Math. \textbf{12} (1961), 334--340. \MR{0132220 (24 \#A2066)}

\bibitem{key-11}
C.-Y. Chen, Y.~You, F.-L. Lu, D.-S. Sun, and S.-H. Dong, \emph{The origin and
  mathematical characteristics of the super-universal associated-{L}egendre
  polynomials}, Commun. Theor. Phys. (Beijing) \textbf{62} (2014), no.~3,
  331--337. \MR{3308824}

\bibitem{key-12}
Y.~M. Chen, Y.~N. Sun, L.~Q. Liu, and X.~H. Ke, \emph{Based on fractional-order
  {L}egendre functions solving a class of fractional-order differential
  equations}, Math. Numer. Sin. \textbf{37} (2015), no.~1, 21--33. \MR{3380905}

\bibitem{key-13}
S.~Choi, J.~Chung, and K.~S. Kim, \emph{Relation between primes and nontrivial
  zeros in the {R}iemann hypothesis; {L}egendre polynomials, modified zeta
  function and {S}chr\"odinger equation}, J. Math. Phys. \textbf{53} (2012),
  no.~12, 122108, 16. \MR{3058172}

\bibitem{key-17}
R.~Courant and D.~Hilbert, \emph{Methods of mathematical physics. {V}ol. {I}},
  Interscience Publishers, Inc., New York, N.Y., 1953. \MR{0065391 (16,426a)}

\bibitem{key-18}
G.~Dattoli, H.~M. Srivastava, and C.~Cesarano, \emph{The {L}aguerre and
  {L}egendre polynomials from an operational point of view}, Appl. Math.
  Comput. \textbf{124} (2001), no.~1, 117--127. \MR{1852063 (2002i:33008)}

\bibitem{key-19}
M.~Hazewinkel, \emph{Legendre polynomials}, Encyclopedia of Mathematics,
  Springer, 2001.

\bibitem{key-20}
I.~R. Horng and J.~H. Chou, \emph{Design of optimal observers with specified
  eigenvalues via shifted {L}egendre polynomials}, J. Optim. Theory Appl.
  \textbf{51} (1986), no.~1, 179--188. \MR{1552019}

\bibitem{key-21}
J.~D. Jackson, \emph{Classical electrodynamics}, second ed., John Wiley \&
  Sons, Inc., New York-London-Sydney, 1975. \MR{0436782 (55 \#9721)}

\bibitem{key-22}
H.~Khalil, R.~A. Khan, M.~H. Al-Smadi, A.~A. Freihat, and N.~Shawagfeh,
  \emph{New operational matrix of shifted {L}egendre polynomials and fractional
  differential equations with variable coefficients}, Punjab Univ. J. Math.
  (Lahore) \textbf{47} (2015), no.~1, 81--103. \MR{3386377}

\bibitem{key-24}
D.~S. Kim and T.~Kim, \emph{Some identities for {B}ernoulli numbers of the
  second kind arising from a non-linear differential equation}, B. Korean Math.
  Soc. \textbf{52} (2015), no.~6, 2001--2010.

\bibitem{key-23}
D.~S. Kim, S.-H. Rim, and T.~Kim, \emph{Some identities on {B}ernoulli and
  {E}uler polynomials arising from orthogonality of {L}egendre polynomials}, J.
  Inequal. Appl. (2012), 2012:227, 8. \MR{3017135}

\bibitem{key-26}
T.~Kim, \emph{Identities involving {L}aguerre polynomials derived from umbral
  calculus}, Russ. J. Math. Phys. \textbf{21} (2014), no.~1, 36--45.
  \MR{3182545}

\bibitem{key-25}
\bysame, \emph{Degenerate {E}uler zeta function}, Russ. J. Math. Phys.
  \textbf{22} (2015), no.~4, 469--472. \MR{3431170}

\bibitem{key-27}
T.~H. Koornwinder, R.~Wong, R.~Koekoek, and R.~F. Swarttouw, \emph{Orthogonal
  polynomials}, N{IST} handbook of mathematical functions, U.S. Dept. Commerce,
  Washington, DC, 2010, pp.~435--484. \MR{2655358}

\bibitem{key-28}
D.~V. Kurdomonov, \emph{An estimate for the rate of convergence of
  {F}ourier-{L}egendre series of functions of bounded variation}, Izv. Vyssh.
  Uchebn. Zaved. Mat. (2006), no.~7, 34--45. \MR{2281887 (2007k:42082)}

\bibitem{key-29}
J.~Kwon and J.-W. Park, \emph{A note on {$(h,q)$}-{B}oole polynomials}, Adv.
  Difference Equ. (2015), 2015:198, 11. \MR{3361795}

\bibitem{key-30}
P.~Mokhtary, F.~Ghoreishi, and H.~M. Srivastava, \emph{The {M}\"untz-{L}egendre
  {T}au method for fractional differential equations}, Appl. Math. Model.
  \textbf{40} (2016), no.~2, 671--684. \MR{3436434}

\bibitem{key-31}
E.~A. Refaat, \emph{Legendre polynomials and functions}, CreateSpace, 2009.

\bibitem{key-32}
D.~Rostamy and S.~Qasemi, \emph{Discontinuous {P}etrov-{G}alerkin and
  {B}ernstein-{L}egendre polynomials method for solving fractional damped heat-
  and wave-like equations}, J. Comput. Theor. Transp. \textbf{44} (2015),
  no.~1, 1--23. \MR{3404328}

\bibitem{key-33}
M.~H. Stone, \emph{Developments in {L}egendre polynomials}, Ann. of Math. (2)
  \textbf{27} (1926), no.~4, 315--329. \MR{1502735}

\bibitem{key-34}
D.~G. Zill and W.~S. Wright, \emph{Advanced {E}ngineering {M}athematics},
  Jones\&Bartlett Publishers, 2009.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amstext}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage{esint}

\makeatletter
\numberwithin{equation}{section}
\numberwithin{figure}{section}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}

\allowdisplaybreaks
\usepackage[sort&compress,square,numbers]{natbib}

\makeatother

\begin{document}

\title{Differential equations associated with Legendre polynomials}

\author{Taekyun Kim}
\address{Department of Mathematics, Kwangwoon University, Seoul 139-701, Republic
of Korea}
\email{tkkim@kw.ac.kr}

\author{Dae San Kim}
\address{Department of Mathematics, Sogang University, Seoul 121-742, Republic
of Korea}
\email{dskim@sogang.ac.kr}

\begin{abstract}
In this paper, we study non-linear differential equations associated with Legendre
polynomials and their applications. From our study of non-linear differential
equations, we derive some new and explicit identities for Legendre polynomials. 
\end{abstract}

\keywords{Legendre polynomials, higher-order Legendre polynomials, non-linear differential equation}

\subjclass[2010]{05A19, 33C45, 34A34}

\maketitle
\global\long\def\relphantom#1{\mathrel{\phantom{{#1}}}}

\section{Introduction}

The Legendre differential equation is given by 
\begin{equation}
\left(1-x^{2}\right)\frac{d^{2}y}{dx^{2}}-2x\frac{dy}{dx}+n\left(n+1\right)y=0,\quad\left(\text{see \cite{key-2,key-4,key-34}}\right).\label{eq:1}
\end{equation}

The equation (\ref{eq:1}) is equivalent to 
\begin{equation}
\frac{d}{dx}\left[\left(1-x^{2}\right)\frac{dy}{dx}\right]+n\left(n+1\right)y=0,\quad\left(\text{see \cite{key-7,key-21}}\right).\label{eq:2}
\end{equation}

The Legendre polynomials (or Legendre functions) are defined as the
solutions of Legendre differential equation. 

The generating function of Legendre polynomials $p_{n}\left(x\right)$
is given by 
\begin{equation}
\frac{1}{\sqrt{1-2xt+t^{2}}}=\sum_{n=0}^{\infty}p_{n}\left(x\right)t^{n},\quad\left(\text{see \cite{key-1,key-2,key-34}}\right).\label{eq:3}
\end{equation}

In physics, the generating function of Legendre polynomials is the
basis for multiple expansion. 

It is known that the Rodrigues' formula of Legendre polynomials is
given by 
\begin{equation}
p_{n}\left(x\right)=\frac{1}{2^{n}n!}\frac{d^{n}}{dx^{n}}\left[\left(x^{2}-1\right)^{n}\right],\quad\left(\text{see \cite{key-34}}\right).\label{eq:4}
\end{equation}

Thus, from (\ref{eq:4}), we note that $p_{n}\left(x\right)$ are
polynomials of degree $n$. 

The Legendre polynomials
$p_{n}\left(x\right)$ can also be represented by the contour integral
as 
\begin{equation}
p_{n}\left(x\right)=\frac{1}{2\pi i}\oint\left(1-2tx+t^{2}\right)^{-\frac{1}{2}}t^{-n-1}dt,\quad\left(\text{see \cite{key-2,key-21}}\right),\label{eq:5}
\end{equation}
where the contour encloses the origin and is traversed in a counterclockwise
direction. 

The first few Legendre polynomials are 
\begin{align*}
p_{0}\left(x\right) & =1,\quad p_{1}\left(x\right)=x,\quad p_{2}\left(x\right)=\frac{1}{2}\left(3x^{2}-1\right),\\
p_{3}\left(x\right) & =\frac{1}{2}\left(5x^{3}-3x\right),\\
p_{4}\left(x\right) & =\frac{1}{8}\left(35x^{4}-30x^{2}+3\right),\quad p_{5}\left(x\right)=\frac{1}{8}\left(63x^{5}-70x^{3}+15x\right),\\
p_{6}\left(x\right) & =\frac{1}{16}\left(231x^{6}-315x^{4}+105x^{2}-5\right),\quad\cdots.
\end{align*}

As is well known, the double factorial of a positive integer $n$
is a generalization of the usual factorial $n!$ defined by Arfken
and given by
\begin{equation}
n!!=\begin{cases}
n\left(n-2\right)\cdots5\cdot3\cdot1 & \text{ if }n>0\text{ odd},\\
n\left(n-2\right)\cdots6\cdot4\cdot2 & \text{if }n>0\text{ even},\\
1 & \text{if }n=-1,0,\quad\left(\text{see \cite{key-21}}\right)
\end{cases}.\label{eq:5-1}
\end{equation}

As was shown by  Arfken, the double factorial can be expressed in terms of the gamma function

\begin{equation}
\Gamma\left(n+\frac{1}{2}\right)=\frac{\left(2n-1\right)!!}{2^{n}}\sqrt{\pi},\quad\left(\text{see \cite{key-8,key-19,key-21}}\right).\label{eq:6}
\end{equation}

Thus, we note that the double factorial can also be extended to negative
odd integers using the definition: 
\begin{equation}
\left(-2n-1\right)!!=\frac{\left(-1\right)^{n}}{\left(2n-1\right)!!}=\frac{\left(-1\right)^{n}2^{n}n!}{\left(2n\right)!}.\label{eq:7}
\end{equation}

Now, we define the higher-order Legendre polynomials as follows: 
\begin{equation}
\left(\frac{1}{\sqrt{1-2tx+t^{2}}}\right)^{\alpha}=\sum_{n=0}^{\infty}p_{n}^{\left(\alpha\right)}\left(x\right)t^{n}.\label{eq:8}
\end{equation}

Some of the explicit formulas for $p_{n}\left(x\right)$ are 
\begin{align*}
p_{n}\left(x\right) & =\frac{1}{2^{n}}\sum_{k=0}^{n}\binom{n}{k}^{2}\left(x-1\right)^{n-k}\left(x+1\right)^{k}\\
 & =\sum_{k=0}^{n}\binom{n}{k}\binom{-n-1}{k}\left(\frac{1-x}{2}\right)^{k}\\
 & =2^{n}\sum_{k=0}^{n}\binom{n}{k}\binom{\frac{n+k-1}{2}}{n},\quad\left(\text{see \cite{key-2,key-34}}\right).
\end{align*}

Legendre polynomials occur in the solution of Laplacian equation of
the static potential $\nabla^{2}\phi\left(x\right)=0$, in a charge-free
region of space, using the method of separation of variables, where
the boundary condition has axial symmetry (no dependence on an azimuthal
angle).

Where $\hat{z}$ is the axis of symmetry and $\theta$ is the angle
between the position of the observer and the $\hat{z}$ axis (the
zenith angle), the solution for the potential will be 
\[
\phi\left(r,\theta\right)=\sum_{l=0}^{\infty}\left[A_{l}r^{l}+B_{l}r^{-\left(l+1\right)}\right]p_{l}\left(\cos\theta\right).
\]

Here $A_{l}$ and $B_{l}$ are to be determined according to the boundary
condition of each problem (see \cite{key-21}). 

Recently, several authors have studied some interesting extensions
and modifications of Legendre polynomials along with related combinatorial,
probabilistic, physics, and physical applications (see \cite{key-1,key-2,key-3,key-4,key-5,key-6,key-7,key-8,key-9,key-10,key-11,key-12,key-13,key-14,key-15,key-16,key-17,key-18,key-19,key-20,key-21,key-22,key-23,key-24,key-25,key-26,key-27,key-28,key-29,key-30,key-31,key-32,key-33}). 

Kim in \cite{key-25,key-26}, and Kim-Kim in \cite{key-24} considered some non-linear differential
equations arising from special numbers and polynomials and derived some new and interesting combinatorial identities. 

In this paper, we consider some differential equations arising from
the generating function of Legendre polynomials and give some new
and explicit identities on the Legendre polynomials which are derived
from the solutions of our differential equations.

\section{Differential equations arising from Legendre polynomials}

Let 
\begin{equation}
F=F\left(t,x\right)=\frac{1}{\sqrt{1-2tx+t^{2}}}.\label{eq:10}
\end{equation}

Then 
\begin{align}
F^{\left(1\right)} & =\frac{d}{dt}F\left(t,x\right)=\left(-\frac{1}{2}\right)\left(1-2tx+t^{2}\right)^{-\frac{3}{2}}\left(-2x+2t\right)\label{eq:11}\\
 & =\left(x-t\right)\left(1-2tx+t^{2}\right)^{-\frac{3}{2}}\nonumber \\
 & =\left(x-t\right)F^{3}.\nonumber 
\end{align}

From (\ref{eq:11}), we have 
\begin{equation}
F^{3}=\frac{1}{x-t}F^{\left(1\right)}.\label{eq:12}
\end{equation}

For $N\in\mathbb{N}$, let 
\begin{equation}
F^{\left(N\right)}=\left(\frac{d}{dt}\right)^{N}F\left(t,x\right),\quad F^{N}=\underset{N-\text{times}}{\underbrace{F\times\cdots\times F}}.\label{eq:13}
\end{equation}

From (\ref{eq:12}), we can derive the following equations: 
\begin{equation}
3F^{2}F^{\left(1\right)}=\frac{\left(-1\right)^{2}}{\left(x-t\right)^{2}}F^{\left(1\right)}+\frac{1}{x-t}F^{\left(2\right)}.\label{eq:14}
\end{equation}

By (\ref{eq:11}) and (\ref{eq:14}), we get 
\begin{equation}
3F^{5}=\frac{1}{\left(x-t\right)^{3}}F^{\left(1\right)}+\frac{1}{\left(x-t\right)^{2}}F^{\left(2\right)}.\label{eq:15}
\end{equation}

From (\ref{eq:15}), we have 
\begin{align}
3\cdot5F^{4}F^{\left(1\right)} & =\frac{3}{\left(x-t\right)^{4}}F^{\left(1\right)}+\frac{1}{\left(x-t\right)^{3}}F^{\left(2\right)}+\frac{2}{\left(x-t\right)^{3}}F^{\left(2\right)}+\frac{1}{\left(x-t\right)^{2}}F^{\left(3\right)}\label{eq:16}\\
 & =\frac{3}{\left(x-t\right)^{4}}F^{\left(1\right)}+\frac{3}{\left(x-t\right)^{3}}F^{\left(2\right)}+\frac{1}{\left(x-t\right)^{2}}F^{\left(3\right)}.\nonumber 
\end{align}

From (\ref{eq:11}) and (\ref{eq:16}), we get 
\begin{equation}
3\cdot5F^{7}=\frac{3}{\left(x-t\right)^{5}}F^{\left(1\right)}+\frac{3}{\left(x-t\right)^{4}}F^{\left(2\right)}+\frac{1}{\left(x-t\right)^{3}}F^{\left(3\right)}.\label{eq:17}
\end{equation}

From (\ref{eq:17}), we note that 
\begin{align}
3\cdot5\cdot7F^{6}F^{\left(1\right)} & =\frac{5\cdot3}{\left(x-t\right)^{6}}F^{\left(1\right)}+\frac{3}{\left(x-t\right)^{5}}F^{\left(2\right)}+\frac{4\cdot3}{\left(x-t\right)^{5}}F^{\left(2\right)}+\frac{3}{\left(x-t\right)^{4}}F^{\left(3\right)}\label{eq:18}\\
 & \relphantom =+\frac{3}{\left(x-t\right)^{4}}F^{\left(3\right)}+\frac{1}{\left(x-t\right)^{3}}F^{\left(4\right)}.\nonumber 
\end{align}

Thus, by (\ref{eq:11}) and (\ref{eq:18}), we get 
\begin{equation}
3\cdot5\cdot7F^{9}=\frac{5\cdot3}{\left(x-t\right)^{7}}F^{\left(1\right)}+\frac{5\cdot3}{\left(x-t\right)^{6}}F^{\left(3\right)}+\frac{3\cdot2}{\left(x-t\right)^{5}}F^{\left(3\right)}+\frac{1}{\left(x-t\right)^{4}}F^{\left(4\right)}.\label{eq:19}
\end{equation}

Continuing this process, we set
\begin{equation}
\left(2N-1\right)!!F^{2N+1}=\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i\right)}}{\left(x-t\right)^{2N-i}}.\label{eq:20}
\end{equation}

Thus, by (\ref{eq:20}), we get 
\begin{align}
 & \left(2N-1\right)!!\left(2N+1\right)F^{2N}F^{\left(1\right)}\label{eq:21}\\
 & =\sum_{i=1}^{N}a_{i}\left(N\right)\frac{d}{dt}\left(\frac{F^{\left(i\right)}}{\left(x-t\right)^{2N-i}}\right)\nonumber \\
 & =\sum_{i=1}^{N}a_{i}\left(N\right)\left\{ \frac{2N-i}{\left(x-t\right)^{2N-i+1}}F^{\left(i\right)}+\frac{1}{\left(x-t\right)^{2N-i}}F^{\left(i+1\right)}\right\} \nonumber \\
 & =\sum_{i=1}^{N}a_{i}\left(N\right)\frac{2N-i}{\left(x-t\right)^{2N-i+1}}F^{\left(i\right)}+\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i+1\right)}}{\left(x-t\right)^{2N-i}}.\nonumber 
\end{align}

From (\ref{eq:11}) and (\ref{eq:21}), we have 
\begin{align}
 & \left(2N+1\right)!!F^{2N+3}\label{eq:22}\\
 & =\sum_{i=1}^{N}a_{i}\left(N\right)\frac{2N-i}{\left(x-t\right)^{2\left(N+1\right)-i}}F^{\left(i\right)}+\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i+1\right)}}{\left(x-t\right)^{2N-i+1}}\nonumber \\
 & =\sum_{i=1}^{N}a_{i}\left(N\right)\frac{2N-i}{\left(x-t\right)^{2\left(N+1\right)-i}}F^{\left(i\right)}+\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i+1\right)}}{\left(x-t\right)^{2N-i+1}}\nonumber \\
 & =\left(2N-1\right)a_{1}\left(N\right)\frac{F^{\left(1\right)}}{\left(x-t\right)^{2N+1}}+a_{N}\left(N\right)\frac{F^{\left(N+1\right)}}{\left(x-t\right)^{N+1}}\nonumber \\
 & \relphantom =+\sum_{i=2}^{N}\left\{ \left(2N-i\right)a_{i}\left(N\right)+a_{i-1}\left(N\right)\right\} \frac{F^{\left(i\right)}}{\left(x-t\right)^{2\left(N+1\right)-i}}.\nonumber 
\end{align}

By replacing $N$ by $N+1$ in (\ref{eq:20}), we get 
\begin{equation}
\left(2N+1\right)!!F^{2N+3}=\sum_{i=1}^{N+1}a_{i}\left(N+1\right)\frac{F^{\left(i\right)}}{\left(x-t\right)^{2\left(N+1\right)-i}}.\label{eq:23}
\end{equation}

By comparing the coefficients on the both sides of (\ref{eq:22})
and (\ref{eq:23}), we have 
\begin{align}
a_{1}\left(N+1\right) & =\left(2N-1\right)a_{1}\left(N\right),\label{eq:24}\\
a_{N+1}\left(N+1\right) & =a_{N}\left(N\right),\label{eq:25}
\end{align}
and 
\begin{equation}
a_{i}\left(N+1\right)=\left(2N-i\right)a_{i}\left(N\right)+a_{i-1}\left(N\right),\quad\left(2\le i\le N\right).\label{eq:26}
\end{equation}

From (\ref{eq:12}) and (\ref{eq:20}), we note that 
\begin{equation}
\frac{1}{x-t}F^{\left(1\right)}=F^{3}=a_{1}\left(1\right)\frac{1}{x-t}F^{\left(1\right)}.\label{eq:27}
\end{equation}

Thus, by comparing the coefficients on both sides of (\ref{eq:27}),
we get $a_{1}\left(1\right)=1$. 

From (\ref{eq:24}) and (\ref{eq:25}), we can derive the following
equations: 
\begin{align}
a_{1}\left(N+1\right) & =\left(2N-1\right)a_{1}\left(N\right)\label{eq:28}\\
 & =\left(2N-1\right)\left(2N-3\right)a_{1}\left(N-1\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(2N-1\right)\left(2N-3\right)\cdots3\cdot1a_{1}\left(1\right)\nonumber \\
 & =\left(2N-1\right)!!,\nonumber 
\end{align}
and 
\begin{equation}
a_{N+1}\left(N+1\right)=a_{N}\left(N\right)=a_{N-1}\left(N-1\right)=\cdots=a_{1}\left(1\right)=1.\label{eq:29}
\end{equation}

By (\ref{eq:26}), for $2\le i \le N$, we get 
\begin{align}
 & a_{i}\left(N+1\right)\label{eq:30}\\
 & =\left(2N-i\right)a_{i}\left(N\right)+a_{i-1}\left(N\right)\nonumber \\
 & =\left(2N-i\right)\left\{ \left(2\left(N-1\right)-i\right)a_{i}\left(N-1\right)+a_{i-1}\left(N-1\right)\right\} +a_{i-1}\left(N\right)\nonumber \\
 & =\left(2N-i\right)\left(2N-2-i\right)a_{i}\left(N-1\right)+\left(2N-i\right)a_{i-1}\left(N-1\right)+a_{i-1}\left(N\right)\nonumber \\
 & =\left(2N-i\right)\left(2N-2-i\right)\left(2N-4-i\right)a_{i}\left(N-2\right)+\left(2N-i\right)\left(2N-2-i\right)a_{i-1}\left(N-2\right)\nonumber \\
 & \relphantom =+\left(2N-i\right)a_{i-1}\left(N-1\right)+a_{i-1}\left(N\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(\prod_{l=0}^{N-i}\left(2N-2l-i\right)\right)a_{i}\left(i\right)+\sum_{l=0}^{N-i}\left\langle 2N-i\right\rangle _{l}a_{i-1}\left(N-l\right)\nonumber \\
 & =\prod_{l=0}^{N-i}\left(2N-2l-i\right)+\sum_{l=0}^{N-i}\left\langle 2N-i\right\rangle _{l}a_{i-1}\left(N-l\right),\nonumber 
\end{align}
where $\left\langle 2N+\alpha\right\rangle _{k}=\left(2N+\alpha\right)\left(2\left(N-1\right)+\alpha\right)\cdots\left(2\left(N-k+1\right)+\alpha\right),$
and $\left\langle 2N+\alpha\right\rangle _{0}=1$. 

From (\ref{eq:30}), we have 
\begin{align}
 & a_{i-1}\left(N-l_{1}\right)\label{eq:31}\\
 & =\prod_{l_{2}=0}^{N-l_{1}-i}\left(2N-2l_{1}-2l_{2}-i-1\right)\nonumber \\
 & \relphantom =+\sum_{l_{2}=0}^{N-l_{1}-i}\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}a_{i-2}\left(N-l_{1}-l_{2}-1\right).\nonumber 
\end{align}

By (\ref{eq:30}) and (\ref{eq:31}), we get 
\begin{align}
 & a_{i}\left(N+1\right)\label{eq:32}\\
 & =\prod_{l_{1}=0}^{N-i}\left(2N-2l_{1}-i\right)+\sum_{l_{1}=0}^{N-i}\left\langle 2N-i\right\rangle _{l_{1}}a_{i-1}\left(N-l_{1}\right)\nonumber \\
 & =\prod_{l_{1}=0}^{N-i}\left(2N-2l_{1}-i\right)+\sum_{l_{1}=0}^{N-i}\left(\prod_{l_{2}=0}^{N-l_{1}-i}\left(2N-2l_{1}-2l_{2}-i-1\right)\right)\left\langle 2N-i\right\rangle _{l_{1}}\nonumber \\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\left\langle 2N-i\right\rangle _{l_{1}}\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}a_{i-2}\left(N-l_{1}-l_{2}-1\right).\nonumber 
\end{align}

Now, we observe that 
\begin{align}
 & a_{i-2}\left(N-l_{1}-l_{2}-1\right)\label{eq:33}\\
 & =a_{i-2}\left(N-l_{1}-l_{2}-2+1\right)\nonumber \\
 & =\prod_{l_{3}=0}^{N-l_{1}-l_{2}-i}\left(2N-2l_{1}-2l_{2}-2l_{3}-i+2\right)\nonumber \\
 & \relphantom =+\sum_{l_{3}=0}^{N-l_{1}-l_{2}-i}\left\langle 2N-2l_{1}-2l_{2}-i-2\right\rangle _{l_{3}}a_{i-3}\left(N-l_{1}-l_{2}-l_{3}-2\right).\nonumber 
\end{align}

From (\ref{eq:32}) and (\ref{eq:33}), and continuing this process,  we obtain
\begin{align*}
 & \relphantom =a_{i}\left(N+1\right)\\
 & =\prod_{l_{1}=0}^{N-i}\left(2N-2l_{1}-i\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\left\langle 2N-i\right\rangle _{l_{1}}\left(\prod_{l_{2}=0}^{N-l_{1}-i}\left(2N-2l_{1}-2l_{2}-i-1\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\left\langle 2N-i\right\rangle _{l_{1}}\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\\
 & \relphantom =\times\left(\prod_{l_{3}=0}^{N-l_{1}-l_{2}-i}\left(2N-2l_{1}-2l_{2}-2l_{3}-i-2\right)\right)+\cdots\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\cdots\sum_{l_{i-2}=0}^{N-l_{1}-\cdots-l_{i-3}-i}\left\langle 2N-i\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-\cdots-2l_{i-3}-2i+3\right\rangle _{l_{i-2}}\\
 & \relphantom =\times\left(\prod_{l_{i-1}=0}^{N-l_{1}-\cdots-l_{i-2}-i}\left(2N-2l_{1}-2l_{2}-\cdots-2l_{i-1}-2i+2\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\cdots\sum_{l_{i-1}=0}^{N-l_{1}-l_{2}-\cdots-l_{i-2}-i}\left\langle 2N-i\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-2l_{2}-\cdots-2l_{i-2}-2i+2\right\rangle _{l_{i-1}}\\
 & \relphantom =\times a_{1}\left(N-l_{1}-l_{2}-\cdots-l_{i-1}-i+2\right)\nonumber \\
 & =\prod_{l_{1}=0}^{N-i}\left(2N-2l_{1}-i\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\left\langle 2N-i\right\rangle _{l_{1}}\left(\prod_{l_{2}=0}^{N-l_{1}-i}\left(2N-2l_{1}-2l_{2}-i-1\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\left\langle 2N-i\right\rangle _{l_{1}}\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\\
 & \relphantom =\times\left(\prod_{l_{3}=0}^{N-l_{1}-l_{2}-i}\left(2N-2l_{1}-2l_{2}-2l_{3}-i-2\right)\right)+\cdots\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\cdots\sum_{l_{i-2}=0}^{N-l_{1}-\cdots-l_{i-3}-i}\left\langle 2N-i\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-\cdots-2l_{i-3}-2i+3\right\rangle _{l_{i-2}}\\
 & \relphantom =\times\left(\prod_{l_{i-1}=0}^{N-l_{1}-\cdots-l_{i-2}-i}\left(2N-2l_{1}-2l_{2}-\cdots-2l_{i-1}-2i+2\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i}\sum_{l_{2}=0}^{N-l_{1}-i}\cdots\sum_{l_{i-1}=0}^{N-l_{1}-l_{2}-\cdots-l_{i-2}-i}\left\langle 2N-i\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-1\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-2l_{2}-\cdots-2l_{i-2}-2i+2\right\rangle _{l_{i-1}}\\
 & \relphantom =\times\left(2\left(N-l_{1}-l_{2}-\cdots-l_{i-1}-i\right)+1\right)!!.
\end{align*}

Therefore, we obtain the following theorem.
\begin{thm}
\label{thm:1} The following non-linear differential equations
\[
\left(2N-1\right)!!F^{2N+1}=\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i\right)}}{\left(x-t\right)^{2N-i}},\quad\left(N=1,2,\dots\right),
\]
has a solution 
\[
F=F\left(t,x\right)=\frac{1}{\sqrt{1-2tx+t^{2}}},
\]
where $a_{1}\left(N\right)=\left(2N-3\right)!!$, $a_{N}\left(N\right)=1$, and, for $2 \le i \le N-1$,
\begin{align*}
 & \relphantom =a_{i}\left(N\right)\\
 & =\prod_{l_{1}=0}^{N-1-i}\left(2N-2l_{1}-i-2\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i-1}\left\langle 2N-2-i\right\rangle _{l_{1}}\left(\prod_{l_{2}=0}^{N-l_{1}-i-1}\left(2N-2l_{1}-2l_{2}-i-3\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i-1}\sum_{l_{2}=0}^{N-l_{1}-i-1}\left\langle 2N-i-2\right\rangle _{l_{1}}\left\langle 2N-2l_{1}-i-3\right\rangle _{l_{2}}\\
 & \relphantom =\times\left(\prod_{l_{3}=0}^{N-l_{1}-l_{2}-i-1}\left(2N-2l_{1}-2l_{2}-2l_{3}-i-4\right)\right)+\cdots\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i-1}\sum_{l_{2}=0}^{N-l_{1}-i-1}\cdots\sum_{l_{i-2}=0}^{N-l_{1}-\cdots-l_{i-3}-i-1}\left\langle 2N-i-2\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-3\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-\cdots-2l_{i-3}-2i+1\right\rangle _{l_{i-2}}\\
 & \relphantom =\times\left(\prod_{l_{i-1}=0}^{N-l_{1}-\cdots-l_{i-2}-i-1}\left(2N-2l_{1}-2l_{2}-\cdots-2l_{i-1}-2i\right)\right)\\
 & \relphantom =+\sum_{l_{1}=0}^{N-i-1}\sum_{l_{2}=0}^{N-l_{1}-i-1}\cdots\sum_{l_{i-1}=0}^{N-l_{1}-l_{2}-\cdots-l_{i-2}-i-1}\left\langle 2N-i-2\right\rangle _{l_{1}}\\
 & \relphantom =\times\left\langle 2N-2l_{1}-i-3\right\rangle _{l_{2}}\cdots\left\langle 2N-2l_{1}-2l_{2}-\cdots-2l_{i-2}-2i\right\rangle _{l_{i-1}}\\
 & \relphantom =\times\left(2\left(N-l_{1}-l_{2}-\cdots-l_{i-1}-i\right)-1\right)!!.
\end{align*}
\end{thm}
Recall that the generating function of Legendre polynomials is given
by 
\begin{equation}
F=F\left(t,x\right)=\frac{1}{\sqrt{1-2tx+t^{2}}}=\sum_{n=0}^{\infty}p_{n}\left(x\right)t^{n}.\label{eq:36}
\end{equation}

Thus, by (\ref{eq:36}), we get 
\begin{align}
F^{\left(i\right)} & =\left(\frac{d}{dt}\right)^{i}F\left(t,x\right)=\sum_{n=i}^{\infty}p_{n}\left(x\right)\left(n\right)_{i}t^{n-i}\label{eq:37}\\
 & =\sum_{n=0}^{\infty}p_{n+i}\left(x\right)\left(n+i\right)_{i}t^{n},\nonumber 
\end{align}
where $\left(x\right)_{n}=x\left(x-1\right)\cdots\left(x-n+1\right)$,
$\left(n\ge1\right)$, and $\left(x\right)_{0}=1$. 

From (\ref{eq:8}), we note that 
\begin{equation}
F^{2N+1}=\sum_{n=0}^{\infty}p_{n}^{\left(2N+1\right)}\left(x\right)t^{n}.\label{eq:38}
\end{equation}

By Theorem \ref{thm:1}, we get 
\begin{align}
 & \relphantom =F^{2N+1}\label{eq:39}\\
 & =\frac{1}{\left(2N-1\right)!!}\sum_{i=1}^{N}a_{i}\left(N\right)\frac{F^{\left(i\right)}}{\left(x-t\right)^{2N-i}}\nonumber \\
 & =\frac{1}{\left(2N-1\right)!!}\sum_{i=1}^{N}a_{i}\left(N\right)\left(\sum_{m=0}^{\infty}\binom{2N+m-i-1}{m}x^{-\left(2N+m-i\right)}t^{m}\right)\nonumber \\
 & \relphantom =\times\left(\sum_{l=0}^{\infty}p_{l+i}\left(x\right)\left(l+i\right)_{i}t^{l}\right)\nonumber \\
 & =\frac{1}{\left(2N-1\right)!!}\sum_{i=1}^{N}a_{i}\left(N\right)\nonumber \\
 & \relphantom =\times\sum_{n=0}^{\infty}\left\{ \sum_{m=0}^{n}\binom{2N+m-i-1}{m}x^{-\left(2N+m-i\right)}p_{n-m+i}\left(x\right)\left(n-m+i\right)_{i}\right\} t^{n}.\nonumber 
\end{align}

Therefore, by (\ref{eq:38}) and (\ref{eq:39}), we obtain the following
theorem.
\begin{thm}
\label{thm:2} For $n\ge0$, we have 
\begin{align*}
p_{n}^{\left(2N+1\right)}\left(x\right) & =\frac{1}{\left(2N-1\right)!!}\sum_{i=1}^{N}\sum_{m=0}^{n}a_{i}\left(N\right)\binom{2N+m-i-1}{m}\\
 & \relphantom =\times x^{-\left(2N+m-i\right)}p_{n-m+i}\left(x\right)\left(n-m+i\right)_{i},
\end{align*}
where $a_{i}\left(N\right)$ $\left(1\le i \le N\right)$ are as in Theorem 1.
\end{thm}
\bibliographystyle{amsplain}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}